Exercise 1-11 Cost Behavior; Contribution Format Income Statement [LO1-4, LO1-6]

Harris Company manufactures and sells a single product. A partially completed schedule of the company’s total costs and costs per unit over the relevant range of 61,000 to 101,000 units is given below:

Required:

1. Complete the schedule of the company’s total costs and costs per unit as given in the relevant tab below.

2. Assume that the company produces and sells 91,000 units during the year at a selling price of $8.37 per unit. Prepare a contribution format income statement for the year.

Units Produced and Sold:

61,000 81,000 101,000

VC: $170,800  

FC: $370,000

Total Costs: $540,800
Answer #1

1)Variable cost Remain same in per unit irrespective of unit produced and sold whereas Fixed cost remain same in Total within a relevant range.

Unit produced and sold
61000 81000 101000
Total cost
variable cost 170800 226800     [2.8*81000] 282,800     [2.8*101000]
Fixed cost 370000 370000 370000
Total cost 540800 596800 652800
costper unit
Variable cost 170800/61000=2.8 2.8 2.8
Fixed cost 370000/61000=6.07 370000/81000=4.57 370000/101000=3.66
Total cost per unit 8.87 7.37 6.46

2)

Contribution Format Income statement
Sales revenue   [91000*8.37] 761,670
variable cost [91000*2.8] (254800)
Contribution margin 506870
less : Fixed cost (370000)
Net Income 136870